Two identical loudspeakers separated by distance  demit 170 Hz sound waves along the x-axis. As you walk along the axis, away from the speakers, you don't hear anything even though both speakers are on.

What are the three lowest possible values for  d? Assume a sound speed of 340 m/s .
Answer #1

Solution,

Wavelength, lambda = 340/170 = 2m

Condition for destructive interference,

d = (m + 0.5) lambda

for the first minima m =0, ±1,± 2 ...

m = 0,

d = lambda /2 = 2/ 2 = 1 m (Ans)

m = 1,

d = 3 lambda /2 = 3 m (Ans)

m = 2,

d = 5 lambda /2 = 5 m (Ans)

Comment in case any doubt please rate my answer ....